OpenStudy (anonymous):

the sum of a certain number and four times the same number is 45. find the number

OpenStudy (anonymous):

Let x be the number you want: x+4x=45 5x=45 x=9

OpenStudy (anonymous):

thank you!!! how bout the sum of two numbers is 38. one of the numbers is ten more than the other. find the larger number.

OpenStudy (anonymous):

x+(x+10) = 38 and solve for x. That will get you the smaller number. Add ten to it to get the larger number.

OpenStudy (anonymous):

so 14 plus 10 is 24

OpenStudy (anonymous):

and 14+24=38 so those sound like the numbers you are looking for!